OpenStudy (anonymous):

3tanx3-10tanx=0 , in the domain 180

OpenStudy (anonymous):

excuse me is it tan3x or tanx3

OpenStudy (anonymous):

tanx^3

OpenStudy (anonymous):

sorry i mean tanx^2

OpenStudy (anonymous):

so 2tanx^2 -10tanx

OpenStudy (anonymous):

omggg 3tanx^2 -10tanx

OpenStudy (anonymous):

@nincompoop

OpenStudy (anonymous):

@abb0t

jimthompson5910 (jim_thompson5910):

Is it \[\Large 3\tan(x^2) - 10\tan(x)=0\] or Is it \[\Large 3\left(\tan(x)\right)^2 - 10\tan(x)=0\] ??

OpenStudy (anonymous):

the second one

jimthompson5910 (jim_thompson5910):

you can let z = tan(x) that gives \[\Large 3z^2 - 10z = 0\] \[\Large z(3z - 10) = 0\] agreed?

OpenStudy (anonymous):

actually the x is the angle sign

OpenStudy (anonymous):

yep i get those steps dk what to di next

jimthompson5910 (jim_thompson5910):

so if z(3z-10) = 0, then either z = 0 or 3z-10 = 0 by the zero product property

jimthompson5910 (jim_thompson5910):

re-substitute in z = tan(x) z = 0 turns into tan(x) = 0 3z-10 = 0 turns into 3*tan(x) - 10 = 0

OpenStudy (anonymous):

jim we need to consider the xrange of x i also got here but the range got me confused

jimthompson5910 (jim_thompson5910):

true, that's a good point

OpenStudy (anonymous):

ok i get this whats next

jimthompson5910 (jim_thompson5910):

solve tan(x) = 0 and 3*tan(x) - 10 = 0 and make sure 180<x<360

jimthompson5910 (jim_thompson5910):

use the arctan function on a calculator to undo the tan function

OpenStudy (anonymous):

or we can put the range

OpenStudy (anonymous):

tan 0 i get 0 , for other i get 73.300

jimthompson5910 (jim_thompson5910):

is x = 0 in the interval 180<x<360 ?

jimthompson5910 (jim_thompson5910):

is x = 73.3 in the interval 180<x<360 ?

OpenStudy (anonymous):

umm

OpenStudy (anonymous):

yea that the part i am stuck on

jimthompson5910 (jim_thompson5910):

if it is, then you can keep the solution if not, then add/subtract 180 until you get in the interval

jimthompson5910 (jim_thompson5910):

being in the interval means that the number is between 180 and 360

OpenStudy (anonymous):

ok 73 is

jimthompson5910 (jim_thompson5910):

nope, 73 is not between 180 and 360

OpenStudy (anonymous):

no i mean when i add 180

jimthompson5910 (jim_thompson5910):

oh ok

OpenStudy (anonymous):

excuse since its periodic the range can be taken in 0<x<180 or 0<x<2pie

OpenStudy (anonymous):

k got the answer thank you

jimthompson5910 (jim_thompson5910):

no problem

OpenStudy (anonymous):

thanks jim

OpenStudy (anonymous):

u guy can stick around ? i got more

OpenStudy (anonymous):

oky woul love to in any way